4x+1 / 3 = x-5 / 2
Answer by deborabr(181)

4x+1 / 3 = x-5 / 2 first make the denominators equal
2(4x+1)/6 = 3(x-5)/6 now delete them
2(4x+1) = 3(x-5)
8x+2 = 3x-15
8x-3x = -15-2
5x = -17
x = -17/5